OpenStudy (anonymous):

what is inverse fourier transform of x(t)= e^-mod t......for -1 < t < 1 = 0 ....otheerwise...

OpenStudy (anonymous):

sorry not inverse .....fourier transform of above function...

hartnn (hartnn):

we know x(t) = e^-t for t=0 to 1 =e^t for t=-1 to 0 so the fourier transform will be split into 2 integrals: \[F(w)=\int\limits_{-1}^{0}e^te^{-jwt}dt+\int\limits_{0}^{1}e^{-t}e^{-jwt}dt\] did u try solving these two integrals ??

OpenStudy (anonymous):

actually i was not getting how to take that mod t.........

OpenStudy (anonymous):

thanks....i got it now.....integration is very easy ....

hartnn (hartnn):

welcome :)